OpenStudy (anonymous):

The slope of a function f at any point (x,y) is y/(2x^2). The point (2,1) is on the graph of f. Find the tangent line to the grpah f at x=2.

OpenStudy (anonymous):

You need a function... what is the function?

OpenStudy (anonymous):

What is\[ \frac{y}{2x^2} \]supposed to be?

OpenStudy (anonymous):

Sorry, didn't type it out word for word my first time around.

OpenStudy (anonymous):

Fixed it now.

OpenStudy (anonymous):

First find the slope: \[ f(x,y) = y/(2x^2) \\ m =f(2,1) = ? \\ \]Here is the equation for a tangent line at \((x_0,y_0)\) with slope \(m\)\[ y-y_0=m(x-x_0) \]